# Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure

### v1

@@ Acknowledgements
We warmly acknowledge Douglas Finkbeiner for very fruitful conversations. A.P. is grateful for the continued support of the DIPC graduate student program as well as the Center for Astrophysics | Harvard & Smithsonian for their warm hospitality at the beginning of this project. R.E. acknowledges the support from grant numbers 21-atp21-0077, NSF AST-1816420, and HST-GO-16173.001-A as well as the Institute for Theory and Computation at the Center for Astrophysics. We are grateful to the supercomputer facility at Harvard University where most of the simulation work was done. G.S. is grateful to the IAS at HKUST for their generous support. This work has been supported by the Spanish project PID2020-114035GB-100 (MINECO/AEI/FEDER, UE). P.M. acknowledges this work was in part performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, Lawrence Livermore National Security, LLC.
Data Availability
@@ Appendix A Stellar Morphology at
Here we present the equivalent plots for data at from the main text, for . This supplementary analysis further enriches our understanding of the core-halo structure in the system under investigation. It is valuable to explore how these parameters interact and correlate at different redshifts, providing a more comprehensive view of the underlying dynamics.

### v2

@@ Acknowledgements.
We warmly acknowledge Douglas Finkbeiner for very fruitful conversations. A.P. is grateful for the continued support of the DIPC graduate student program as well as the Center for Astrophysics — Harvard & Smithsonian for their warm hospitality at the beginning of this project. R.E. acknowledges the support from grant numbers 21-atp21-0077, NSF AST-1816420, and HST-GO-16173.001-A as well as the Institute for Theory and Computation at the Center for Astrophysics. We are grateful to the supercomputer facility at Harvard University where most of the simulation work was done. G.S. is grateful to the IAS at HKUST for their generous support. This work has been supported by the Spanish project PID2020-114035GB-100 (MINECO/AEI/FEDER, UE). P.M. acknowledges this work was in part performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, Lawrence Livermore National Security, LLC.Data availability
@@ Appendix A Stellar Morphology at
Here we present the equivalent plots for data at from the main text, for . This supplementary analysis further enriches our understanding of the core-halo structure in the system under investigation. It is valuable to explore how these parameters interact and correlate at different redshifts, providing a more comprehensive view of the underlying dynamics. Figures 10, 11 and 12.
Appendix B Tables with metallicity, age and spatial distribution of stars
Here we present a summary in two tables ( Tables 3 and 4) of the metallicities, ages, and spatial distribution of the stars in the three simulated halos. These tables are interesting for quantitatively observing the distribution of stars in terms of age, metallicity, and spatial location within both the cores and halos of the simulated Galaxies across the three dark matter models.
| Halos | (-2¡[Fe/H]¡0) | (-4¡[Fe/H]¡-2) | (-6¡[Fe/H]¡-4) | (-8¡[Fe/H]¡-6) | |
| DMG1 | 12340 | 93.12 | 4.91 | 0.71 | 1.26 |
| DMG2 | 1950 | 80.62 | 14.82 | 1.44 | 3.12 |
| DMG3 | 1550 | 77.22 | 16.91 | 1.68 | 4.19 |
| CDMG1 | 21206 | 95.62 | 3.61 | 0.36 | 0.41 |
| CDMG2 | 7572 | 88.55 | 9.17 | 1.12 | 1.16 |
| CDMG3 | 3417 | 83.79 | 12.38 | 1.82 | 2.01 |
| “WDM”G1 | 28328 | 89.48 | 8.43 | 0.76 | 1.33 |
| “WDM”G2 | 3793 | 79.25 | 14.79 | 2.14 | 3.82 |
| “WDM”G3 | 5122 | 78.21 | 16.34 | 1.44 | 4.01 |
| Halos | ||||||||
| DMG1 | 12340 | 11.57 | 8.02 | 0.92 | 48.72 | 37.41 | 10.46 | 3.41 |
| DMG1 | 1950 | 5.66 | 2.82 | 0.16 | 72.10 | 16.41 | 7.23 | 4.26 |
| DMG1 | 1550 | 5.10 | 2.89 | 0.15 | 64.45 | 27.36 | 8.19 | 0 |
| CDMG1 | 21206 | 13.21 | 5.58 | 0.74 | 48.60 | 37.27 | 7.75 | 6.38 |
| CDMG2 | 7572 | 24.57 | 8.24 | 2.02 | 29.32 | 30.78 | 14.05 | 25.85 |
| CDMG3 | 3417 | 7.99 | 10.87 | 0.87 | 35.91 | 24.00 | 9.51 | 30.58 |
| “WDM”G1 | 28328 | 24.56 | 13.75 | 3.37 | 38.04 | 36.55 | 12.05 | 13.36 |
| “WDM”G2 | 3793 | 10.18 | 6.14 | 0.63 | 63.88 | 20.43 | 6.52 | 9.17 |
| “WDM”G3 | 5122 | 16.08 | 9.31 | 1.50 | 49.57 | 38.54 | 11.21 | 0.68 |
Appendix C Corner Plots from MCMC fit
Here we present the resulting corner plot of the performed MCMC fit corresponding to Figure 6. Figure 13.